Using the z-table, find the corresponding area between 𝑧 = 0 and each of the following:

1. z = 0.92

2. z = 1.29

3. z = 2.73

4. z = βˆ’0.50

5. z = βˆ’2.98



1
Expert's answer
2021-04-15T07:20:18-0400

1. "\Phi(0.92)-\Phi(0)=0.82121-0.5"

=0.32121

2. "\Phi (1.29)-\Phi (0)=0.90147-0.5"

=0.40147

3. "\Phi (2.73)-\Phi (0)=0.99683-0.5"

=0.49673

4. "\Phi (0)-\Phi (-0.5)= 0.5-(1-0.69146)"

=0.19146

5. "\Phi (0)-\Phi (-2.98)=0.5-(1-0.99856)"

=0.49856